OpenStudy (anonymous):

Find the smallest possible value of \( 2x^2 - 3x + 4 \)

OpenStudy (anonymous):

find the derivative then set it equal to zero

OpenStudy (jamesj):

Or complete the square and read off the minimum.

OpenStudy (anonymous):

4x - 3 = 0 x = 3/4 then plug this in the equation

OpenStudy (hoblos):

21/8

OpenStudy (anonymous):

2(x^2 - 3/2 x + 9/4) + 4 - 9/4 = 2(x - 3/2)^2 + 4 - 9/4 is it 4 - 9/4 ?

OpenStudy (anonymous):

I think I did a tiny mistake

OpenStudy (anonymous):

you need to subtract 2(9/4) from the outside of the parenthesis to balance the equation.

OpenStudy (anonymous):

okay... Ill try again

OpenStudy (anonymous):

you have everything else right, it just needs to be:\[4-\frac{18}{4}=-\frac{1}{2}\]

OpenStudy (anonymous):

2x^2 - 3x + 4 = 2 (x^2 - 3/2 x + 2) = 2(x^2 - 3/2 x) + 4 why am I even struggling with this :-)

OpenStudy (anonymous):

but the answer book says the smallest value is actually 23/8

OpenStudy (anonymous):

it should be 2(x - 3/4)^2 + 4 - 9/8

OpenStudy (anonymous):

oh I see my mistake now

OpenStudy (anonymous):

'and at that moment, Agdgdgdgwngo was enlightened'

OpenStudy (anonymous):

yeah i just saw it too >.< there should be a 9/16's inside.

OpenStudy (jamesj):

Yes, 2x^2−3x+4 = 2(x^2 - 3/2.x +2) = 2[ (x - 3/4)^2 + 23/16 ]